# Sensor Placement via Large Deviations in the Eikonal Equation

### v1

@@ 1. Introduction
These problems can then be formulated in a shape optimization framework. Indeed, given a set $\Omega\subset\mathbb{R}^{2}$, and a mass fraction $c\in(0,|\Omega|)$, we state the following problem
@@
By employing a homogenization strategy, in which the mass of many micro-sensors is uniformly distributed throughout $\Omega$, one can show that Problem (1) does not admit an actual solution. Indeed, the infimum of the cost functional vanishes and is asymptotically approached by a sequence of disconnected sets with an increasing number of connected components densely covering $\Omega$. However, there is clearly no sensor configuration $\omega$ for which the maximal distance becomes zero, unless $c=|\Omega|$.
@@
The main contribution of this paper is the proposal to employ a classical result of Varadhan [22] to approximate the distance function through the solution of a suitable elliptic PDE. This approach naturally leads to a family of approximated problems (see Section 3), for which we perform a sensitivity analysis using shape-derivative techniques (see Section 4.1). These tools are then efficiently applied to numerical optimization, yielding the computational results presented in Section 4.
@@ 2. State of the art and Applications
A first example arises in gas networks, where strategically positioning sensors is crucial for the efficient distribution and management of resources such as natural gas and hydrogen. Sensors are typically deployed at critical locations?pipelines, compressor stations, and distribution hubs?to monitor key parameters including pressure, temperature, flow rates, and gas composition. Such deployments enable the rapid detection of leaks or pressure fluctuations, allowing operators to respond promptly and thereby mitigate safety risks and environmental damage. Moreover, the collected data supports performance optimization, reduces energy losses, and enhances the overall reliability and cost-effectiveness of the network.
@@
  • The problem of designing actuators within a prescribed set, to optimize the control performance of a process, has been extensively investigated in the framework of optimal control. Several methodologies for effective actuator placement have been developed, notably those based on spectral techniques, as discussed in [18, 19, 20], with applications primarily to control problems for wave and heat processes.
@@
  • This perspective has been applied to mathematical biology problems, such as optimizing resource distribution to maximize population size [16, 17], as well as the mathematical analysis of optimal habitat configurations for species persistence [21].
@@
It is important to note, however, that in this paper we approach the problem from a purely geometric perspective, without reference to any specific application. In this sense, the sensor location problems we address have a form of universal validity, as they rely solely on geometric considerations. The outcomes of our methodologies can thus be viewed as an initialization step, to be further refined when incorporating the specific knowledge of a given application–for instance, by constraining the admissible functions to be solutions of a particular partial differential equation, as is often the practice case.
@@ 3. Approximating the distance function via a Varadhan’s result
Figure 1. The box $D$ containing the domaine $\Omega$ and the sensors $(B_{i})$.
@@
As we will see below, although efficient and straightforward to implement, gradient methods suffer from a fundamental limitation since they only guarantee convergence to a local minimum, not necessarily to the global one.
Developing efficient methods that guarantee convergence to the global minimizer is a challenging problem and lies beyond the scope of this paper.
@@ 4.1. Computation of the gradients
$\begin{cases}u^{\prime}-\varepsilon\Delta u^{\prime}=-0&\qquad\mbox{in $D\backslash\omega$,}\vskip 2.84526pt\\ \ \ \ \ \ \ \ \ \ \ u^{\prime}=-\frac{\partial u}{\partial\nu}\langle V,\nu\rangle d\sigma&\qquad\mbox{on $\partial\omega\cup\partial D$.}\end{cases}$
It is classical that
@@
$\int_{D\backslash\omega}q\varphi dx-\varepsilon\int_{D\backslash\omega}\Delta q\varphi dx=-\int_{D\backslash\omega}j^{\prime}(u)u^{\prime}\mathbbm{1}_{\Omega}dx=-\int_{\Omega\backslash\omega}j^{\prime}(u)u^{\prime}\varphi dx.$
@@ 4.2. Numerical experiments
$p=1$ $p=5$ $p=10$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
$p=1$ $p=2$ $p=10$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
$p=1$ $p=2$ $p=10$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ 5. Conclusion and perspectives
Acknowledgement. I. Ftouhi was supported by the Alexander von Humboldt Foundation via a Postdoctoral fellowship.

### v2

@@ 1. Introduction
These problems can then be formulated in a shape optimization framework. Indeed, given a set $\Omega\subset\mathbb{R}^{n}$, and a mass fraction $c\in(0,|\Omega|)$, we state the following problem
@@
By employing a homogenization strategy, in which the mass of many micro-sensors is uniformly distributed throughout $\Omega$, one can show that Problem (1) does not admit an actual solution. Indeed, the infimum of the cost functional vanishes and is asymptotically approached by a sequence of disconnected sets with an increasing number of connected components densely covering $\Omega$. Note that, in the absence of additional geometric constraints on the sensor set $\omega$, one can always construct sensors of zero measure whose Hausdorff distance to $\Omega$ is zero, i.e., $d^{H}(\omega,\Omega)=0$. For instance, taking $\omega=\Omega\cap\mathbb{Q}^{d}$ yields a set of Lebesgue measure zero that is dense in $\Omega$, and hence satisfies $d^{H}(\omega,\Omega)=0$. By considering sufficiently small neighborhoods around arbitrarily distributed rational points, this construction can be extended to sets of any prescribed measure.
@@
The main contribution of this paper is the use of a classical result of Varadhan [27], stated in Theorem 3.1, to approximate the distance function through the solution of a suitable elliptic PDE. This approach naturally leads to a family of approximated problems (see Section 3 and more precisely Lemma 3.2), for which we perform a sensitivity analysis using shape-derivative techniques, see Section 4.1 where a shape derivation result is presented in Theorem 4.1. Then, these tools are applied efficiently to numerical optimization, producing the numerical results presented in Section 4.
@@ 2. State of the art and Applications
A first example arises in gas networks, where strategically positioning sensors is crucial for the efficient distribution and management of resources such as natural gas and hydrogen. Sensors are typically deployed at critical locations, such as pipelines, compressor stations, and distribution hubs to monitor key parameters including pressure, temperature, flow rates, and gas composition. Such deployments enable the rapid detection of leaks or pressure fluctuations, allowing operators to respond promptly and thereby mitigate safety risks and environmental damage. Moreover, the collected data supports performance optimization, reduces energy losses, and enhances the overall reliability and cost-effectiveness of the network.
@@
  • The problem of designing actuators within a prescribed set, to optimize the control performance of a process, has been extensively investigated in the framework of optimal control. Several methodologies for effective actuator placement have been developed, notably those based on spectral techniques, as discussed in [22, 23, 24], with applications primarily to control problems for wave and heat processes. We also refer to [1, 17, 19, 28] for a non-exhaustive list of references linking theory with applications.
@@
  • This perspective has been applied to mathematical biology problems, such as optimizing resource distribution to maximize population size [20, 21], as well as the mathematical analysis of optimal habitat configurations for species persistence [26]. Also, optimal sensor placement plays a crucial role in the study of the brain, as it directly determines the quality, interpretability, and efficiency of the data acquired from neuroimaging and electrophysiological techniques such as electroencephalogram (EEG) and near-infrared spectroscopy (NIRS), see for example [7, 25] .
@@
It is important to note, however, that in this paper we approach the problem from a purely geometric perspective, without reference to any specific application. In this sense, the sensor location problems we address have a form of universal validity, as they rely solely on geometric considerations. The outcomes of our methodologies can thus be viewed as an initialization step, to be further refined when incorporating the specific knowledge of a given application–for instance, by constraining the admissible functions to be solutions of a particular partial differential equation, as is often the case in practice.
@@ 3. Approximating the distance function via a Varadhan’s result
Figure 1. The box $D$ containing the domain $\Omega$ and the sensors $(B_{i})$. As it can be seen in the figure, the box $D$ is chosen such that $d(B_{4},\partial\Omega)=d(B_{4},M_{4})\leq d(M_{4},\partial D)$
@@
As we will see in remark 4.4, although efficient and straightforward to implement, gradient methods suffer from a fundamental limitation since they only guarantee convergence to a critical point or a local minimum, when a second order optimality condition is used, not necessarily to the global one. Developing efficient methods that guarantee convergence to the global minimizer is a challenging problem and lies beyond the scope of this paper.
@@ 4.1. Computation of the gradients
$\begin{cases}u^{\prime}-\varepsilon\Delta u^{\prime}=-0&\qquad\mbox{in $D\backslash\omega$,}\vskip 2.84526pt\\ \ \ \ \ \ \ \ \ \ \ u^{\prime}=-\frac{\partial u}{\partial\nu}\langle V,\nu\rangle&\qquad\mbox{on $\partial\omega\cup\partial D$.}\end{cases}$
The last claims are standard, we refer to [15, Section 5.3] for the formal proofs. On the other hand, we have
@@
$\int_{D\backslash\omega}q\varphi dx-\varepsilon\int_{D\backslash\omega}\Delta q\varphi dx=-\int_{D\backslash\omega}j^{\prime}(u)\varphi\mathbbm{1}_{\Omega}dx=-\int_{\Omega\backslash\omega}j^{\prime}(u)\varphi dx.$
@@ 4.2. Numerical experiments
The simulations of both methods were conducted using Matlab software on a personal DELL computer with a quadcore Intel Core i7-2600 with 16GB RAM. The resolution of PDE (6) was done by using finite element method via the toolbox PDEtool and with the parameter $\varepsilon$ taken equal to $10^{-3}$. The optimization procedure was performed through ’fmincon’ routine with the following settings:
  • Algorithm: "sqp".
  • ConstraintTolerance: 1.0000e-06.
  • OptimalityTolerance: 1.0000e-06.
  • StepTolerance: 1.0000e-06.
  • SpecifyObjectiveGradient: "on".
@@
Remark 4.3.
The optimal placement of the sensors is not necessarily unique. For example, if we consider the domain $\Omega$ to be a disk and take $N\geq 2$ sensors, then any rotation of an optimal placement with respect to the center of (the disk) $\Omega$ is also optimal.
Remark 4.4.
As explained above, the use of gradient descent can lead to critical points in our case since we only use a first order sensitivity analysis. For example, if we consider two sensors, $\Omega$ to be a square and $p=1$, we notice that the method will converge to different critical points (with different values of the mean distances) depending on the choice of the initial placement of the sensors as shown in Figure 5 below:
Refer to caption Refer to caption
Mean distance $\approx 0.8066$ Mean distance $\approx 0.7947$
Figure 5. Two critical configurations of the sensors obtained for $p=1$ by considering different initial positions: $(0.3,0)$ and (0.3,0)(-0.3,0) for the case on the left and $(0.1,0.1)$ and (0.3,0.1)(-0.3,-0.1) for the case on the right, which seems to be a global maximum.
@@ 5. Conclusion and perspectives
Acknowledgement. The work of I. Ftouhi was supported by the Alexander von Humboldt Foundation via a Postdoctoral fellowship and also partially supported by the French National Research Agency under the France 2030 program, grant ANR-23-EXES-0005 (project Gardener/MASRA), and by the Occitanie Region.